OpenStudy (anonymous):

Which of the following is the reciprocal of x^2 – 2x – 8? A. (x + 2)(x – 4) B. 1/ (x + 2)(x – 4) C. 1/ (x – 8) D. 1/(x + 2)

OpenStudy (anonymous):

The area of a rectangle is x^2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x – 5 B. x^2 + 15 C. x + 5 D. (x – 2)(x + 3)

OpenStudy (anonymous):

I need help with both questions please?

OpenStudy (mathteacher1729):

What does the word "reciprocal" mean?

OpenStudy (anonymous):

i got the answer to the first one.

OpenStudy (anonymous):

need help with the second one

OpenStudy (anonymous):

The area of a rectangle is x^2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x – 5 B. x^2 + 15 C. x + 5 D. (x – 2)(x + 3)

OpenStudy (anonymous):

the answer to the first is b. 1/(x+2)(x-4)

OpenStudy (mathteacher1729):

Area of a rectangle is Length * Width ==> A = L*W L = (x+3) A = (x+3) * W x^2 - 2x - 15 = (x+3) * W So can you factor x^2 - 2x - 15 = (x+3)( ) ? The missing piece is the answer.

OpenStudy (anonymous):

x-2

OpenStudy (mathteacher1729):

Are you saying that (x+3)(x+2) = x^2 -2x -15 ?

OpenStudy (anonymous):

no i'm saying (x+3)(x-2)

OpenStudy (mathteacher1729):

Oh sorry. (x+3)(x-2) = x^2 -2x -15 Check the constant terms... 3*-2 = -6 not -15

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3*-5=-15 so it is x-5

OpenStudy (anonymous):

What is the remainder of (x^3 – 6x – 9x + 3) ÷ (x - 3) A. -51 B. -51/x - 3 C. -17/x D. -17/ x - 3

OpenStudy (mathteacher1729):

Hey I'm leaving now. Time to get back to work for me. Good luck with the rest of your problems.

OpenStudy (anonymous):

can you help me with this last one